OpenStudy (anonymous):

Solve the system of equations 4x + 3y = 13 y=- x + 4

pooja195 (pooja195):

ok which method would you like to use?

OpenStudy (anonymous):

I totally get these backwards, could you help figure this out

pooja195 (pooja195):

\[4x+3(-x+4)=13\] the x has a negative right and lets use substitution here its easier

pooja195 (pooja195):

all you do if take y=- x + 4 and plug it in place of the 1st equation

pooja195 (pooja195):

now solve from there

OpenStudy (anonymous):

do I add or subtract from the 13

pooja195 (pooja195):

no first distribut 3 so 3 x -x = ? 3 x4 =?

OpenStudy (anonymous):

7 x - 13

pooja195 (pooja195):

no....we are MULTIPLYING 3 x -x = ? 3 x4 =?

OpenStudy (anonymous):

ok then 12 x

pooja195 (pooja195):

the x was for multiplication not 12x just 12 whats 3 times -x ?

pooja195 (pooja195):

\[4x+3(-x+4)=13\] distribute \[4x-3x+12=13\] combine like terms \[1x=1\] divide x=1

OpenStudy (anonymous):

-12 right divided by 12 = 1

pooja195 (pooja195):

I already did that part for you now we have to substitute for X on the 2nd equation \[y=-(1)+4\] substitute 1 in place of X \[y=-1+4\] \[y=3\]

pooja195 (pooja195):

so your final answer is?

OpenStudy (anonymous):

(3, 4) would this be the answer?

pooja195 (pooja195):

where did u get 4 from?

pooja195 (pooja195):

i already did all the work now u make the coordinates i tell u what equals what x= ? y= ?

OpenStudy (anonymous):

(3,12)

pooja195 (pooja195):

Look back at my work

pooja195 (pooja195):

if x=1 y=3 what is your coordinate?

OpenStudy (anonymous):

-3,-16 ? I am confused

pooja195 (pooja195):

I gave u the values there right there \[X=1~~~Y=3\]

pooja195 (pooja195):

Now make the coordinates

OpenStudy (anonymous):

ok, now I see it would be (1,3)

pooja195 (pooja195):

\(\huge\color{RED}{CONGRATS!!!!!}\)

OpenStudy (anonymous):

Thanks so much, now i am going to try to figure this out online. I will let you know how I did in a bit.

pooja195 (pooja195):

your welcome =)

OpenStudy (anonymous):

I have to go somewhere soon, so thanks again pooja

pooja195 (pooja195):

hahah your welcome :) also welcome to openstudy! :D

OpenStudy (anonymous):

thanks, there is so much more to learn, I forgot how to do all of these problems :)

pooja195 (pooja195):

hehe tag me anytime :)

OpenStudy (domebotnos):

@wopwop if anythings you now know that pooja is always ready for action and the rest of us owls :) welcome :D